Kiana has made 9 of 19 freethrows so far this season. Her goal is to make 60% of her freethrows. If Kiana makes her next x freethrows in a row, the function f(x) = (9 + x)/ (19 + x) represents Kiana’s new ratio of freethrows made. How many successful freethrows in a row will raise Kiana’s percent made to 60% ?
Answer by richwmiller(17219)

(9+x)/(19+x)=3/5
x=6
if she makes the next 6 freethrows she'll have 60%
15/25=3/5= 60%